OpenStudy (anonymous):

using a scale of this map is 1 inch = 30 miles. how many actual miles does 5.25 inches represent a)2.2 inch x 3.3 inch b.)2.4 inch x 3.6 inch c.)2.8 inch x 4.2inch d.)2.9inch x 5.0 inch

OpenStudy (mathstudent55):

The choices are for a different problem.

OpenStudy (anonymous):

oh yea xD

OpenStudy (mathstudent55):

1 inch = 30 miles 2 inches = 2 * 30 miles 3 inches = 3 * 30 miles 4 inches = 4 * 30 miles Do you see a pattern? Now do 5.25 inches = and fill in the right side.

OpenStudy (anonymous):

im not smart so i dont understand a thing you did ._.

OpenStudy (anonymous):

but the answer choices are a.)157.5 miles b.)175 miles c.)1,575 miles d.) 1,70 miles

OpenStudy (anonymous):

its B 175

OpenStudy (anonymous):

is it a ?

OpenStudy (anonymous):

how o_0

OpenStudy (anonymous):

ive been sick for the past week . and i dont have a clue how to do this

OpenStudy (anonymous):

or yeah sorry its a i divided instead of multiply sorry

OpenStudy (anonymous):

Get better! :)

OpenStudy (mathstudent55):

You are given a scale. 1 inch on the map means 30 miles of real distance. Are you following so far?

OpenStudy (anonymous):

so its a ?

OpenStudy (anonymous):

Thanks !

OpenStudy (anonymous):

im trying to figure out the answer it is really confusing me so is it a geesh -..-

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

I'm trying to explain it to you. I can continue if you want me to.

OpenStudy (anonymous):

No Problem!

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

i jsut need to know if my answer is right because i am timed on this test which id ont have no idea about whats its on

OpenStudy (anonymous):

5.25 * 30 = 157.5

OpenStudy (mathstudent55):

1 inch on the map means 30 miles of real distance. If you multiply the statement above by 2, you get: 2 inches on the map means 60 miles of real distance. Ok so far?